OpenStudy (anonymous):

inverse equation of f(x)=3x+4 please show all steps

jimthompson5910 (jim_thompson5910):

The idea is to replace f(x), swap x and y, then solve for y. f(x)=3x+4 y=3x+4 x=3y+4 x-4=3y (x-4)/3 = y y = (x-4)/3 So the inverse function is \[\Large f^{-1}(x) = \frac{x-4}{3}\]

jimthompson5910 (jim_thompson5910):

oops meant to write "replace f(x) with y"

OpenStudy (anonymous):

wow thanks a lot, exact answer i was looking for can u please stand by me to help I have a couple more i need help with can I talk to you on aim or something?

jimthompson5910 (jim_thompson5910):

you can talk to me through here, msn messenger or yahoo messenger

jimthompson5910 (jim_thompson5910):

yahoo is probably best for me

OpenStudy (anonymous):

okay cool can I add you on msn so i can keep in touch with you whenever i need help with another problem?

OpenStudy (anonymous):

i mean yahoo**

jimthompson5910 (jim_thompson5910):

sure you can do that